An ionic compound A2B dissociates in water according to the equation

A2B(s) ⇌ 2A+(aq) + B2–(aq)

If the concentration of B2– ions in a saturated solution of A2B is 3.2 × 10–3 M, then what is the solubility product constant (Ksp) of this compound?

  •  A. 8.2 × 10–7
  •  B. 3.3 × 10–8
  •  C. 1.3 × 10–7
  •  D. 5.6 × 10–7
  •  E. 2.2 × 10–7
1
Expert's answer
2021-03-10T06:15:14-0500

Solution:

An ionic compound A2B dissociates in water according to the equation:

A2B(s) ⇌ 2A+(aq) + B2–(aq)


[B2–] = 3.2×10–3 M

[A+] = 2 × [B2–] = 2 × 3.2×10–3 M = 6.4×10–3 M (according to the equation above)

Ksp = [A+]2 × [B2–]

Hence,

Ksp = (6.4×10–3)2 × (3.2×10–3) = 1.31×10–7 = 1.3×10–7

Ksp = 1.3×10–7


Answer: C. 1.3×10–7